# Motion Polynomials Admitting a Factorization with Linear Factors

### v3

@@ 3.5 Factorizations of General Bounded Motion Polynomials
The final step is to extend the result of Proposition 4 to motion polynomials with arbitrary norm polynomial. The idea is to combine Proposition 3, Lemma 3, and a technique called “Bennett flips” introduced in [20]. It is based on replacing a factorization $(t-h_{1})(t-h_{2})$ of a generic, monic motion polynomial of degree two by its second factorization $(t-k_{1})(t-k_{2})$ [9, Theorem 1]. Note that $\nu(t-h_{1})=\nu(t-k_{2})$ and $\nu(t-h_{2})=\nu(t-k_{1})$. First, we generalize the latter procedure.
@@
Consider a bounded monic reduced motion polynomial $M=c+\varepsilon D\in\mathbb{D}\mathbb{H}[t]$ of degree $n$ with $c\in\mathbb{R}[t]$ and a linear bounded monic reduced motion polynomial $t-p-\varepsilon d\in\mathbb{D}\mathbb{H}[t]$ with $p,q\in\mathbb{H}$. Let $\nu(M)=c^{2}=N_{1}^{n}$ and $\nu(t-p-\varepsilon d)=N_{2}$ for some irreducible, quadratic polynomials $N_{1}$, $N_{2}\in\mathbb{R}[t]$ with $\operatorname{realgcd}(N_{1},N_{2})=1$. If $M$ admits a factorization into monic linear motion polynomial factors, then there exists a quaternion polynomial $D_{1}\in\mathbb{H}[t]$ and a quaternion $d_{1}\in\mathbb{H}$ such that $M(t-p-\varepsilon d)=(t-p-\varepsilon d_{1})(c+\varepsilon D_{1})$, and $c+\varepsilon D_{1}$ admits a factorization into monic linear motion polynomial factors.
@@
Consider a bounded, monic, reduced motion polynomial $M=cQ+\varepsilon D\in\mathbb{DH}[t]$ with $c\in\mathbb{R}[t]$ and $\operatorname{realgcd}(Q)=1$. Let $M=\prod_{i=1}^{r}(c_{i}Q_{i}+\varepsilon D_{i})$ be a factorization of $M$ given by Proposition 3 with $c_{i}\in\mathbb{R}[t]$ and $\operatorname{realgcd}(Q_{i})=1$. For $i=1,\ldots,r$, set
@@
Consider a bounded, monic, reduced motion polynomial $M=cQ+\varepsilon D\in\mathbb{DH}[t]$ with $c\in\mathbb{R}[t]$ and $\operatorname{realgcd}(Q)=1$. Define $g\coloneqq\operatorname{realgcd}(c,D{Q}^{\ast},{Q}^{\ast}D)$. If $M$ admits a factorization into monic linear factors, then $cg$ divides $\nu(D)$.
@@
Consider a bounded, monic, reduced motion polynomial $M=cQ+\varepsilon D\in\mathbb{DH}[t]$ with $c\in\mathbb{R}[t]$ and $\operatorname{realgcd}(Q)=1$. Set $g\coloneqq\operatorname{realgcd}(c,{Q}^{\ast}D,D{Q}^{\ast})\in\mathbb{R}[t]$. If $cg$ divides $\nu(D)$, then $M$ admits a factorization into monic linear factors.
@@ 3.6 A Comprehensive Example
  • Since $\nu(M)=(t^{2}+1)^{4}$, $c=\operatorname{realgcd}P=t^{2}+1$, $Q=(t-\mathbf{i})^{2}$, $D=\mathbf{i}(t-\mathbf{i})^{2}$, $g=\operatorname{realgcd}(c,{Q}^{\ast}D,{D}^{\ast}Q)=t^{2}+1$, and $cg$ divides $\nu(D)$, we can apply Algorithm 4.
@@ 4 Another proof of Theorem 1
where $cg|\nu(D)$ and $\nu(D)=N\nu(D_{1})$. It remains to prove that $\tau(c_{1}g_{1})<\tau(cg)$. By the assumption $\tau(D{P}^{\ast})\geq\tau({P}^{\ast}D)=\tau({P_{1}}^{\ast}D_{1})+1$, we get
@@
Since $c^{2}|{P}^{\ast}P,{P}^{\ast}vP,(vP-Pv){(vP-Pv)}^{\ast}$ and $g|c^{2},D{P}^{\ast},{P}^{\ast}D$, it follows that the transformation preserves the condition $g|D{D}^{\ast}$ and thus the assumptions of the theorem.
@@
1:A bounded monic reduced motion polynomial $M=P+\varepsilon D\in\mathbb{D}\mathbb{H}[t]$ such that $\operatorname{realgcd}(\operatorname{realgcd}(P)^{2},{P}^{\ast}D,D{P}^{\ast})|{D}^{\ast}D$.
@@
3:If $\deg\operatorname{realgcd}(P)=0$ Then
@@
6:     $N\leftarrow$ any quadratic real factor of $\operatorname{realgcd}(P)$
@@
17:              $q\leftarrow$ any nonzero vectorial quaternion such that ${p}^{\ast}q=qp$
18:              $M_{1}\leftarrow P_{1}+\varepsilon(D_{1}+qP/N)$
19:              Return $[t-p-\varepsilon q,\texttt{Factor4}(M_{1})]$
20:         End If
@@
Consider a bounded, monic, reduced motion polynomial $M=cQ+\varepsilon D\in\mathbb{DH}[t]$ with $c\in\mathbb{R}[t]$ and $\operatorname{realgcd}(Q)=1$. Set $g\coloneqq\operatorname{realgcd}(c,{Q}^{\ast}D,D{Q}^{\ast})\in\mathbb{R}[t]$ and $g^{\prime}\coloneqq\frac{cg}{\operatorname{realgcd}(cg,\nu{(D)})}\in\mathbb{R}[t]$. Then $Mg^{\prime}$ admits a factorization into monic linear factors.
@@
To perform the induction step, take a bounded monic reduced motion polynomial $M_{1}$ such that $\deg g_{1}^{\prime}\neq 0$ Here $g_{1}^{\prime},g_{1},c_{1},D_{1},Q_{1}$ for $M_{1}$ are defined analogously to $g^{\prime},g,c,D,Q$. Pick a monic real quadratic irreducible factor $N$ of $g^{\prime}_{1}$. Assume without loss of generality that $\tau({Q_{1}}^{\ast}D_{1})\leq\tau(D_{1}{Q_{1}}^{\ast})$, because $M_{1}g_{1}^{\prime}$ admits a factorization if and only if ${M_{1}}^{\ast}g_{1}^{\prime}$ does, and the conjugation of $M_{1}$ interchanges ${Q_{1}}^{\ast}D_{1}$ and $D_{1}{Q_{1}}^{\ast}$. Divide the ratio ${Q_{1}}^{\ast}D_{1}/\operatorname{realgcd}({Q_{1}}^{\ast}D_{1})$ by $N$ with a remainder $qt+r$. Since $M_{1}$ is a motion polynomial, it follows that ${Q_{1}}^{\ast}D_{1}+{D_{1}}^{\ast}Q_{1}=0$, hence $q$ and $r$ are vectorial quaternions not vanishing simultaneously. There are infinitely many quaternions $p$ such that $\nu(t-p)=N$ and $p(pq+r)\neq(pq+r)p$. Indeed, consider the vectorial quaternion $r^{\prime}:=r+q(p+{p}^{\ast})/2$ determined by $q$, $r$, and $\nu(t-p)=N$. Take the vectorial part of $p$ to be orthogonal to $q$ if $r^{\prime}=0$ or $q$ is non-orthogonal to $r^{\prime}$, orthogonal to $r^{\prime}$ if $q=0$, and orthogonal to neither $q$ nor $r^{\prime}$ if $q\neq 0$ is orthogonal to $r^{\prime}$. Then $p(pq+r)-(pq+r)p$ is not orthogonal to $r^{\prime}$ (for $r^{\prime}\neq 0$) and hence nonzero. Among the infinitely many $p$, pick one such that $t-{p}^{\ast}$ is not a right factor of $Q_{1}$ and $D_{1}$.
Take $M:=M_{1}(t-p)$. We have $\operatorname{realgcd}({Q}^{\ast}D)=\operatorname{realgcd}({Q_{1}}^{\ast}D_{1})$ because
$\frac{{Q}^{\ast}D}{\operatorname{realgcd}({Q_{1}}^{\ast}D_{1})}$ $=(t-{p}^{\ast})\frac{{Q_{1}}^{\ast}D_{1}}{\operatorname{realgcd}({Q_{1}}^{\ast}D_{1})}(t-p)=(t-{p}^{\ast})(qt+r)(t-p)$

### v4

@@ 3.5 Factorizations of General Bounded Motion Polynomials
The final step is to extend the result of Proposition 4 to motion polynomials with arbitrary norms. We combine Proposition 3, Lemma 3, and a technique called “Bennett flips” introduced in [20]. It is based on replacing a factorization $(t-h_{1})(t-h_{2})$ of a generic monic motion polynomial of degree two with $\nu(t-h_{1})\neq\nu(t-h_{2})$ by its second factorization $(t-k_{1})(t-k_{2})$ [9, Theorem 1]. Note that $\nu(t-h_{1})=\nu(t-k_{2})$ and $\nu(t-h_{2})=\nu(t-k_{1})$. First, we generalize the latter procedure.
@@
Consider a bounded monic reduced motion polynomial $M=c+\varepsilon D\in\mathbb{D}\mathbb{H}[t]$ of degree $n$ with $c\in\mathbb{R}[t]$ and a linear bounded monic reduced motion polynomial $t-p-\varepsilon d\in\mathbb{D}\mathbb{H}[t]$ with $p,d\in\mathbb{H}$. Let $\nu(M)=c^{2}=N_{1}^{n}$ and $\nu(t-p-\varepsilon d)=N_{2}$ for some irreducible, quadratic polynomials $N_{1}$, $N_{2}\in\mathbb{R}[t]$ with $\operatorname{realgcd}(N_{1},N_{2})=1$. If $M$ admits a factorization into monic linear motion polynomial factors, then there exists a quaternion polynomial $D_{1}\in\mathbb{H}[t]$ and a quaternion $d_{1}\in\mathbb{H}$ such that $M(t-p-\varepsilon d)=(t-p-\varepsilon d_{1})(c+\varepsilon D_{1})$, and $c+\varepsilon D_{1}$ admits a factorization into monic linear motion polynomial factors.
@@
Consider a bounded, monic, reduced motion polynomial $M=cQ+\varepsilon D\in\mathbb{DH}[t]$ with $c\in\mathbb{R}[t]$ and $\operatorname{mrpf}(Q)=1$. Let $M=\prod_{i=1}^{r}(c_{i}Q_{i}+\varepsilon D_{i})$ be a factorization of $M$ given by Proposition 3 with $c_{i}\in\mathbb{R}[t]$ and $\operatorname{mrpf}(Q_{i})=1$. For $i=1,\ldots,r$, set
@@
Consider a bounded, monic, reduced motion polynomial $M=cQ+\varepsilon D\in\mathbb{DH}[t]$ with $c\in\mathbb{R}[t]$ and $\operatorname{mrpf}(Q)=1$. Define $g\coloneqq\operatorname{realgcd}(c,D{Q}^{\ast},{Q}^{\ast}D)$. If $M$ admits a factorization into monic linear factors, then $cg$ divides $\nu(D)$.
@@
Consider a bounded, monic, reduced motion polynomial $M=cQ+\varepsilon D\in\mathbb{DH}[t]$ with $c\in\mathbb{R}[t]$ and $\operatorname{mrpf}(Q)=1$. Set $g\coloneqq\operatorname{realgcd}(c,{Q}^{\ast}D,D{Q}^{\ast})\in\mathbb{R}[t]$. If $cg$ divides $\nu(D)$, then $M$ admits a factorization into monic linear factors.
@@ 3.6 A Comprehensive Example
  • Since $\nu(M)=(t^{2}+1)^{4}$, $c=\operatorname{mrpf}P=t^{2}+1$, $Q=(t-\mathbf{i})^{2}$, $D=\mathbf{i}(t-\mathbf{i})^{2}$, $g=\operatorname{realgcd}(c,{Q}^{\ast}D,{D}^{\ast}Q)=t^{2}+1$, and $cg$ divides $\nu(D)$, we can apply Algorithm 4.
@@ 4 Another proof of Theorem 1
where $c_{1}g_{1}|cg|\nu(D)$ and $\nu(D)=N\nu(D_{1})$. It remains to prove that $\tau(c_{1}g_{1})<\tau(cg)$. By the assumption $\tau(D{P}^{\ast})\geq\tau({P}^{\ast}D)=\tau({P_{1}}^{\ast}D_{1})+1$, we get
@@
Since $c^{2}|{P}^{\ast}P,{P}^{\ast}vP,(vP-Pv){(vP-Pv)}^{\ast}$ and $cg|c^{2},D{P}^{\ast},{P}^{\ast}D$, it follows that the transformation preserves $c$, $g$, the condition $cg|D{D}^{\ast}$, and thus the assumptions of the theorem.
@@
1:A bounded monic reduced motion polynomial $M=P+\varepsilon D\in\mathbb{D}\mathbb{H}[t]$ such that $\operatorname{realgcd}(\operatorname{mrpf}(P)^{2},{P}^{\ast}D,D{P}^{\ast})|{D}^{\ast}D$.
@@
3:If $\deg\operatorname{mrpf}(P)=0$ Then
@@
6:     $N\leftarrow$ any quadratic real factor of $\operatorname{mrpf}(P)$
@@
17:              $v\leftarrow$ any vectorial quaternion that does not commute with $p$
18:              $q\leftarrow pv-vp$
19:              $M_{1}\leftarrow P_{1}+\varepsilon(D_{1}+qP/N)$
20:              Return $[t-p-\varepsilon q,\texttt{Factor4}(M_{1})]$
@@
Consider a bounded, monic, reduced motion polynomial $M=cQ+\varepsilon D\in\mathbb{DH}[t]$ with $c\in\mathbb{R}[t]$ and $\operatorname{mrpf}(Q)=1$. Set $g\coloneqq\operatorname{realgcd}(c,{Q}^{\ast}D,D{Q}^{\ast})\in\mathbb{R}[t]$ and $g^{\prime}\coloneqq\frac{cg}{\operatorname{realgcd}(cg,\nu{(D)})}\in\mathbb{R}[t]$. Then $Mg^{\prime}$ admits a factorization into monic linear factors.
@@
To perform the induction step, take a bounded monic reduced motion polynomial $M_{1}$ such that $\deg g_{1}^{\prime}\neq 0$ Here $g_{1}^{\prime},g_{1},c_{1},D_{1},Q_{1}$ for $M_{1}$ are defined analogously to $g^{\prime},g,c,D,Q$. Pick a monic real quadratic irreducible factor $N$ of $g^{\prime}_{1}$. Assume without loss of generality that $\tau({Q_{1}}^{\ast}D_{1})\leq\tau(D_{1}{Q_{1}}^{\ast})$, because $M_{1}g_{1}^{\prime}$ admits a factorization if and only if ${M_{1}}^{\ast}g_{1}^{\prime}$ does, and the conjugation of $M_{1}$ interchanges ${Q_{1}}^{\ast}D_{1}$ and $D_{1}{Q_{1}}^{\ast}$. Divide the ratio ${Q_{1}}^{\ast}D_{1}/\operatorname{mrpf}({Q_{1}}^{\ast}D_{1})$ by $N$ with a remainder $qt+r$. Since $M_{1}$ is a motion polynomial, it follows that ${Q_{1}}^{\ast}D_{1}+{D_{1}}^{\ast}Q_{1}=0$, hence $q$ and $r$ are vectorial quaternions not vanishing simultaneously. There are infinitely many quaternions $p$ such that $\nu(t-p)=N$ and $p(pq+r)\neq(pq+r)p$. Indeed, consider the vectorial quaternion $r^{\prime}:=r+q(p+{p}^{\ast})/2$ determined by $q$, $r$, and $\nu(t-p)=N$. Take the vectorial part of $p$ to be orthogonal to $q$ if $r^{\prime}=0$ or $q$ is non-orthogonal to $r^{\prime}$, orthogonal to $r^{\prime}$ if $q=0$, and orthogonal to neither $q$ nor $r^{\prime}$ if $q\neq 0$ is orthogonal to $r^{\prime}\neq 0$. Then $p(pq+r)-(pq+r)p$ is not orthogonal to $r^{\prime}$ (for $q,r^{\prime}\neq 0$) and hence nonzero. Among the infinitely many $p$, pick one such that $t-{p}^{\ast}$ is not a right factor of $Q_{1}$ and $D_{1}$.
Take $M:=M_{1}(t-p)$. We have $\operatorname{mrpf}({Q}^{\ast}D)=\operatorname{mrpf}({Q_{1}}^{\ast}D_{1})$ because
$\frac{{Q}^{\ast}D}{\operatorname{mrpf}({Q_{1}}^{\ast}D_{1})}$ $=(t-{p}^{\ast})\frac{{Q_{1}}^{\ast}D_{1}}{\operatorname{mrpf}({Q_{1}}^{\ast}D_{1})}(t-p)=(t-{p}^{\ast})(qt+r)(t-p)$
@@ 5 Conclusion
Acknowledgments
Zijia Li is supported by the National Key R&\&&D Program of China (2023YFA1009401) and partially supported by the Strategic Priority Research Program of the Chinese Academy of Sciences $0$ XDB0640200 and the Guangdong Basic and Applied Basic Research Foundation under Grant 2024A1515010506. Mikhail Skopenkov is supported in part by the KAUST baseline. Daniel Scharler was supported by the Austrian Science Fund (FWF) P 33397-N (Rotor Polynomials: Algebra and Geometry of Conformal Motions). The first two authors acknowledge the support of the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université) and LabEx CARMIN (ANR-10-LABX-59-01).